# Active Hypothesis Testing under Computational Budgets with Applications to GWAS and LLM

### v1

@@ 3 Hypothesis Testing under Budget Constraint
The admissibility results in Section 2.4 establish a fundamental property of our framework: statistical power is not derived from a universally optimal control function, but rather from a data-adaptive strategy that intelligently allocates the global budget $n_{b}$ across the $N$ hypotheses. We now return to the problem formulated in Section 2.1 and present such a strategy in detail.
@@ 3.1 A Normalized Allocation Scheme
$h_{i}(\mathbf{X}^{a})=\frac{n_{b}\cdot u_{i}(X_{i}^{a})}{\sum_{j=1}^{N}u_{j}(X_{j}^{a})}.$ (8)
This construction is data-driven, allocating a proportionally larger share of the budget to hypotheses with auxiliary statistics deemed more promising. Furthermore, it is designed to satisfy $\sum_{i=1}^{N}h_{i}(\mathbf{X}^{a})=n_{b}$.
Recall that $C_{i}=\mathbb{I}(\text{statistic }X_{i}\text{ is computed})$. Then the expected total cost is bounded by
$\mathbb{E}\Big[\sum_{i=1}^{N}C_{i}\Big]=\sum_{i=1}^{N}\mathbb{E}[\min(1,h_{i}(\mathbf{X}^{a}))]\leq\sum_{i=1}^{N}\mathbb{E}[h_{i}(\mathbf{X}^{a})]=\mathbb{E}\Big[\sum_{i=1}^{N}h_{i}(\mathbf{X}^{a})\Big]=n_{b}.$
So the scheme rigorously adheres to the budget in expectation.
Beyond the expectation-level guarantee, the actual number of computations concentrates sharply around the budget $n_{b}$, as it is a sum of independent Bernoulli random variables. Specifically, by choosing utility functions such that $h_{i}\leq 1$ (cf. Section 3.2), we can use the concentration inequality to show that for any $\delta>0$ and $\epsilon>0$, the total number of expensive computations lies within the interval $n_{b}-\epsilon,n_{b}+\epsilon]$ with probability at least $1-\delta$, provided that the sample size satisfies $N\geq\frac{2\epsilon^{2}}{\log(2/\delta)}$.
@@ 3.2 Guidance on Selecting the Utility Functions
This strategy creates a powerful synergistic effect. Consider the active $e$-value construction (2). Under the alternative, a promising auxiliary statistic (e.g., a large $E_{i}^{a}$ in the direct signal case) will produce a large utility $u_{i}(E_{i}^{a})$, which in turn increases its control value $h_{i}(E_{i}^{a})$. This has two synergistic benefits:
@@
Practical Refinements.
A crucial practical issue is that the realized number of computations may sometimes fall substantially below the target budget $n_{b}$. To understand why, note that for any given realization of $\mathbf{X}^{a}$, the expected number of computations is $\sum_{i=1}^{N}\min(1,h_{i}(\mathbf{X}^{a}))$, which is always less than or equal to $\sum_{i=1}^{N}h_{i}(\mathbf{X}^{a})=n_{b}$. This inequality becomes strict whenever any control value $h_{i}(\mathbf{X}^{a})>1$. This occurs when the distribution of utility values $\{u_{i}(X_{i}^{a})\}_{i=1}^{N}$ is highly dispersed, with a few outliers being exceptionally large. In such cases, the computational budget cannot be fully utilized.
To address this, we can apply some more range-compressing utility functions to make sure $h_{i}(\mathbf{X}^{a})\leq 1$ for any $i=1,\ldots,N$. A common and effective choice is the logarithm:
$u_{i}(x)=\log(1+x)\quad(\text{for a direct signal}),\quad u_{i}(x)=\log\left(1+\frac{1}{x+\epsilon}\right)\quad(\text{for an inverse signal}).$
This calibration step mitigates the impact of outliers and promotes fuller use of the budget by enabling a more balanced allocation.
@@ 3.3 Budgeted Active Inference Algorithm
With the data-adaptive allocation strategy and active statistic constructions in place, we now present the complete, unified algorithm for budgeted active inference. The procedure, detailed in Algorithm 1, takes as input the auxiliary statistics, a global budget, the hyperparameter $\beta$, and user-specified control functions. It returns a valid test statistic for every hypothesis, having rigorously adhered to the budget constraint.
@@
0: Auxiliary statistics $\{X_{i}^{a}\}_{i=1}^{N}$, global budget $n_{b}$, hyperparameter $\beta\in(0,1)$, control functions $\{h_{i}(\cdot)\}_{i=1}^{N}$.
1:for each hypothesis $i=1,\dots,N$ do
2:  Draw an independent random variable $U_{i}\sim\mathrm{Uniform}(0,1)$.
3:  if $U_{i}<h_{i}$ then
4:   Compute the exact statistic $X_{i}$.
5:   Construct the active statistic based on $X_{i}$:
@@
    • $b_{i}=\frac{h_{i}}{1-\beta}$ (independent),
    • $b_{i}=\frac{\min(1,\sup_{\mathbf{y}\in\mathbb{R}^{N}}h_{i}(\mathbf{y}))}{1-\beta}$ (general).
6:  else
7:   Construct the active statistic without $X_{i}$:
@@
8:  end if
9:end for
10:Return: The set of active statistics $\{X_{i}^{\mathrm{active}}\}_{i=1}^{N}$.
The set of active statistics $\{X_{i}^{\mathrm{active}}\}_{i=1}^{N}$ produced by Algorithm 1 is designed to be broadly compatible with a wide range of downstream multiple testing procedures, a key advantage of our framework. This design allows researchers to choose the procedure that best suits the form of the statistic produced (whether a $p$-value or an $e$-value), the dependence structure in the data, and the desired power for controlling error metrics such as the False Discovery Rate (FDR, Benjamini and Hochberg, (1995)).
For instance, the resulting active $p$-values can be supplied to a spectrum of methods tailored to different dependency assumptions. These range from the classic Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg,, 1995), which is powerful under independence or PRDS, to the Su procedure, which provides guarantees under the PRDN assumption (Su,, 2018), and the highly robust Benjamini-Yekutieli (BY) procedure (Benjamini and Yekutieli,, 2001) for arbitrary dependence. Moreover, they are compatible with more advanced techniques, including adaptive procedures that estimate the null proportion $\pi_{0}$ to boost power (Storey,, 2002; Storey et al.,, 2004) and sophisticated conditional calibration methods like the dBH procedure (Fithian and Lei,, 2022).
Alternatively, when formulated as $e$-values, our statistics can be integrated with modern $e$-value-based methods, which are particularly appealing for their robustness to complex dependencies. Notable examples include the standard e-BH procedure for arbitrary dependence (Wang and Ramdas,, 2022), enhanced methods that boost power via conditional calibration such as e-BH-CC (Lee and Ren,, 2024), and unifying frameworks like the e-Closure Principle introduced by Xu et al., 2025a , which can offer uniform improvements in power and flexibility. Our framework thus serves as a flexible front-end, compatible with this entire suite of modern statistical machinery.
@@ 4 Numerical Experiments
We conduct a series of numerical simulations to evaluate our budgeted active inference framework, comparing its statistical power and efficiency against several baselines under a fixed budget. The utility functions, $\{u_{i}\}_{i=1}^{N}$, are chosen for each experiment as guided by Section 3.2. Unless otherwise specified, we use the identity function $u_{i}(x)=x$ for direct signals and $u_{i}(x)=1/(x+\epsilon)$ with $\epsilon=10^{-8}$ for inverse signals. To handle settings with highly dispersed utilities we may apply the logarithm in the end.
@@ 4.1 Competing Methods and Evaluation Metrics
We compare the proposed method “Active-Default” with the following methods:
@@
    ALL (Oracle). This non-budgeted oracle method computes the exact statistic ($E_{i}$ or $P_{i}$) for all $N$ hypotheses. It serves as an upper bound on statistical power at a fixed cost of $N$ queries.
@@
    Xu (Xu et al., 2025b, ). This method makes an independent probabilistic decision for each hypothesis. A Bernoulli trial $T_{i}\sim\mathrm{Bernoulli}(p_{i})$ determines whether to compute the expensive statistic, where the probability $p_{i}$ is a function of the auxiliary statistic and a hyperparameter $\beta$. For $e$-values, the query probability is $p_{i}=\max\{0,1-\beta/E_{i}^{a}\}$, and the final statistic is $E_{i}^{\mathrm{active}}=(1-T_{i})E_{i}^{a}+T_{i}(1-\beta)E_{i}$. For $p$-values, $p_{i}=\max\{0,1-\beta P_{i}^{a}\}$, and the final statistic is $P_{i}^{\mathrm{active}}=(1-T_{i})P_{i}^{a}+T_{i}(1-\beta)^{-1}P_{i}$. Crucially, because these decisions are made independently for each hypothesis, the total number of queries $\sum_{i}T_{i}$ is a random variable and is not constrained by a pre-specified global budget.
@@ 4.2 Performance with an Auxiliary Signal
In our first experiment, we assess performance in a scenario where the auxiliary statistic provides a direct but unquantifiable signal about the true effect. We simulate $N=10,000$ hypotheses, each defined by a signal strength parameter $\mu_{i}$. The $i$th null hypothesis is $H_{0,i}:\mu_{i}=0$. The signal strengths $\{\mu_{i}\}_{i=1}^{N}$ are generated independently from a two-component mixture model to create a fraction $\pi$ of non-nulls:
@@
Let $\tau^{2}=2\log N$. From each primary observation $Z_{i}\sim\mathcal{N}(\mu_{i},1)$, we construct a corresponding gold-standard $e$-value and $p$-value: $E_{i}=\exp\left(\lambda Z_{i}-\frac{\lambda^{2}}{2}\right)\text{ and }P_{i}=1-\Phi(Z_{i}),$ where $\lambda=\sqrt{\log(N/\alpha)}$ as recommended in Xu et al., 2025b and $\Phi$ is the standard normal CDF. The corresponding auxiliary statistics, which encode the signal strength $\mu_{i}$, are generated as: $E_{i}^{a}\sim\mathrm{Poisson}(1+\mu_{i})\text{ and }P_{i}^{a}\sim\mathrm{Beta}(1,1+\mu_{i}).$ The Poisson statistic serves as a direct signal for the $e$-value, while the Beta statistic provides an inverse signal for the $p$-value.
We conduct two analyses based on this setup, with a computational budget of $n_{b}=500$. First, to assess performance as a function of signal density, we vary the non-null proportion $\pi$ from 0.05 to 0.3 while holding $\beta=0.5$ fixed. Second, to examine the influence of the hyperparameter $\beta$, we vary it from 0.1 to 0.9 while keeping $\pi=0.1$ fixed.
For our proposed Active-Default method, we use the utility function $u_{i}(x)=x$ for the direct signal ($e$-values) and $u_{i}(x)=\log(1+1/(x+\epsilon))$ for the inverse signal ($p$-values). Given that the statistics are generated independently for each hypothesis, we employ the active $p$-value construction designed for the independent case as in (6). The results for each analysis, averaged over 100 simulations, are presented in Figure 1 and Figure 2, respectively.
@@
4.3 Performance with a Noisy Proxy
$E_{i}$ $=\exp\!\left(\lambda Z_{i}-\dfrac{\lambda^{2}}{2}\right)\qquad\text{and}\qquad P_{i}=1-\Phi(Z_{i}),$ (9)
$E_{i}^{a}$ $=\exp\!\left(\lambda Y_{i}-\dfrac{\lambda^{2}}{2}\right)\qquad\text{and}\qquad P_{i}^{a}=1-\Phi(Y_{i}).$ (10)
We conduct two analyses within this framework. First, we fix the noise standard deviation at a moderate level of $\sigma=1$ and vary the non-null proportion $\pi$ from 0.05 to 0.3. Second, we fix $\pi=0.1$ and vary $\sigma$ from 1 to 5 to assess the methods’ robustness to deteriorating proxy quality. Due to the high dispersion of the auxiliary statistics in this setup, we use the range-compressing utility function $u_{i}(x)=\log(1+x)$ for the $e$-value setting and $\log(1+1/(x+\epsilon))$ for the $p$-value setting to ensure the budget was fully utilized. Here we adopt the active $p$-value constructed for the general dependent case as in (7).
Figure 3: Performance comparison as a function of $\pi$, with a fixed $\sigma=1$.
The results of the first analysis, shown in Figure 3, are highly consistent with our findings from Section 4.2. All methods control the FDR, and our globally budgeted approaches perfectly adhere to the $n_{b}=500$ query limit. The Active-Default method again emerges as the most efficient, with its advantage growing as the proportion of true signals increases.
Figure 4: Performance comparison as a function of $\sigma$, with a fixed $\pi=0.1$.
The second analysis, presented in Figure 4, probes the methods’ robustness. As the noise level $\sigma$ increases, the auxiliary statistic $E_{i}^{a}$ becomes a less reliable indicator of the exact $e$-value $E_{i}$. Consequently, the power and efficiency of Xu, Active-Xu, and Active-Default decline. However, the performance ranking remains stable. Our Active-Default method consistently outperforms the other budget-constrained methods across all noise levels. This demonstrates that even as the quality of the auxiliary information degrades, our framework’s ability to efficiently allocate a fixed budget provides a durable performance advantage.
4.4 Performance with a Correlated Proxy
Our third and final simulation investigates a scenario where the gold-standard and auxiliary statistics share a deeper structural relationship modeled by correlation. This is representative of many real-world problems where a cheap measurement is an indirect but correlated indicator of an expensive one (e.g., gene expression levels and protein abundance).
[…]
We then compute $(E_{i},E_{i}^{a})$ and $(P_{i},P_{i}^{a})$ from $(Z_{i},Y_{i})$ via the definitions given in (9) and (10), and the correlation $\rho$ directly controls the quality of both auxiliary channels.
We perform two analyses. First, we fix the correlation at a moderate level, $\rho=0.5$, and vary $\pi$ from 0.05 to 0.3. Second, we fix $\pi=0.1$ and vary $\rho$ from 0.2 to 0.9, assessing how well each method capitalizes on improving proxy quality. As in the default choice, we use $u_{i}(x)=\log(1+x)$ for the $e$-value setting and $u_{i}(x)=\log(1+1/(x+\epsilon))$ for the $p$-value setting. Again, we adopt the active $p$-value constructed for the general dependent case as in (7).
Figure 5: Performance comparison as a function of $\pi$, with a fixed $\rho=0.5$.
The results of the first analysis, displayed in Figure 5, confirm the robustness of our method. The performance patterns are consistent with those observed in the previous, structurally different simulations. Active-Default adheres to the budget while delivering the highest efficiency, with its advantage widening as the proportion of true signals grows.
Figure 6: Performance comparison as a function of $\rho$, with a fixed $\pi=0.1$.
The second analysis, shown in Figure 6, provides a deeper insight into the methods’ behavior. As $\rho$ increases, the auxiliary statistic becomes a more faithful proxy for the gold-standard statistic. This increased information quality allows all active inference methods to improve their power and efficiency. However, Active-Default demonstrates the most significant gains. Its efficiency curve rises more steeply than those of the other methods, highlighting its superior ability to capitalize on high-quality side information. This result shows that our framework not only works well with weak proxies but excels when strong auxiliary data are available, making it an adaptive and powerful tool for budgeted inference.

### v2

@@ 3 Hypothesis Testing under Budget Constraint
The admissibility results in Section 2.4 establish a fundamental property of our framework: statistical power is not derived from a universally optimal control function, but rather from a data-adaptive strategy that intelligently allocates the global budget $n_{b}$ across the NN hypotheses. We now return to the problem formulated in Section 2.1 and present such a strategy.
@@ 3.1 A Normalized Allocation Scheme
$h_{i}(\mathbf{X}^{a})=n_{b}\cdot\frac{u_{i}(X_{i}^{a})}{\sum_{j=1}^{N}u_{j}(X_{j}^{a})}.$ (8)
By construction, this mathematically ensures the exact sum constraint $\sum_{i=1}^{N}h_{i}(\mathbf{X}^{a})=n_{b}$.
@@ 3.2 Guidance on Selecting the Utility Functions
However, if the base utilities are highly skewed, naively computing allocations via the normalized scheme (8) may yield $h_{i}>1$. Simply capping $h_{i}$ at 11 would cause $\sum h_{i}<n_{b}$, and the available resources will not be utilized fully. To guarantee $h_{i}\in[0,1]$, we employ an adaptive transformation applied to the base utilities: $u_{i}(x)=\log(1+(u_{i}^{\text{base}}(x))^{1/k})$, where $k$ is a positive integer. Intuitively, taking the logarithm reduces large differences among the base utilities, and increasing the integer $k$ enforces a progressively stronger compression. Because $n_{b}\leq N$, there always exists an integer $k$ sufficiently large to guarantee $\max_{i}h_{i}\leq 1$. Crucially, this adaptive compression step relies solely on the auxiliary statistics $\{X_{i}^{a}\}$, so it is computationally inexpensive.
This utility selection strategy creates a strong synergy. Consider the active $e$-value construction (2). Under the alternative, a promising auxiliary statistic (e.g., a large $E_{i}^{a}$ in the direct signal case) will produce a large utility $u_{i}(E_{i}^{a})$, which in turn increases its control value $h_{i}(E_{i}^{a})$. This yields two benefits:
@@ 3.3 Budgeted Active Inference Algorithm
Next, a central technical challenge is to ensure that the total number of expensive computations exactly equals $n_{b}$ on every run. Unlike previous methods (Xu et al., 2025b) that rely on independent coin flips which result in random, unpredictable budget utilization, our framework requires a dependent sampling mechanism that correlates the decisions across all NN hypotheses to ensure strict budget adherence. Formally, we seek to sample binary indicators $C_{1},\dots,C_{N}\in\{0,1\}$ conditionally on $\mathbf{X}^{a}$ such that they satisfy two conditions simultaneously: valid marginal selection probabilities, meaning $C_{i}\mid\mathbf{X}^{a}\sim\mathrm{Bernoulli}(h_{i}(\mathbf{X}^{a}))$ for each $i$; and exact global budget adherence, meaning $\sum_{i=1}^{N}C_{i}=n_{b}$. While the theoretical existence of such a joint distribution is guaranteed by Chen et al. (2022), this existence result does not directly yield a practical sampling algorithm. To address this, the next proposition provides an explicit construction of $C_{1},\ldots,C_{N}$ that satisfies these conditions.
Proposition 3.
Suppose $p_{1},\dots,p_{N}\in[0,1]$ satisfy $\sum_{i=1}^{N}p_{i}=n_{b}\in\mathbb{N}$. Let $S_{i}=\sum_{j=1}^{i}p_{j}$ for $i=1,\dots,N$, with $S_{0}=0$ and $U\sim\mathrm{Uniform}(0,1)$. Define
$C_{i}=\lfloor S_{i}-U\rfloor-\lfloor S_{i-1}-U\rfloor,\quad i=1,\dots,N.$ (9)
Then marginally $C_{i}\sim\text{Bernoulli}(p_{i})$ for all $i$, and $\sum_{i=1}^{N}C_{i}=n_{b}$.
We are now ready to present the complete algorithm for budgeted active inference. The procedure, detailed in Algorithm 1, takes as input the auxiliary statistics, a global exact budget, the hyperparameter $\beta$, and user-specified utility functions. It returns a valid test statistic for every hypothesis and rigorously adheres to the constraints.
@@
0: Auxiliary statistics $\{X_{i}^{a}\}_{i=1}^{N}$, exact global budget $n_{b}$, hyperparameter $\beta\in(0,1)$, user-specified base utility functions $\{u_{i}(\cdot)\}_{i=1}^{N}$.
1:Step 1: Normalized Allocation
2: Let $u_{i}=u_{i}(X_{i}^{a})$ for $i=1,\dots,N$ and $k=1$.
3: Set $h_{i}=n_{b}\cdot\frac{u_{i}}{\sum_{j}u_{j}}$ for $i=1,\dots,N$.
4:while $\max_{i}h_{i}>1$ do
5:  Adaptively compress utilities: $u_{i}\leftarrow\log(1+(u_{i})^{1/k})$ for $i=1,\dots,N$.
6:  Recompute $h_{i}=n_{b}\cdot\frac{u_{i}}{\sum_{j}u_{j}}$ for $i=1,\dots,N$.
7:  Update $k\leftarrow k+1$.
8:end while
9:Step 2: Exact-Sum Dependent Sampling
10: Compute cumulative limits $S_{i}=\sum_{j=1}^{i}h_{j}$ (with $S_{0}=0$).
11: Sample $U\sim\mathrm{Uniform}(0,1)$.
12:for each hypothesis $i=1,\dots,N$ do
13:  Set indicator boolean $C_{i}=\lfloor S_{i}-U\rfloor-\lfloor S_{i-1}-U\rfloor$.
14:  if $C_{i}=1$ then
15:   Compute the exact primary statistic $X_{i}$.
16:   Construct the active statistic based on $X_{i}$:
@@
    • $b_{i}=\frac{h_{i}}{1-\beta}$ (under independence, i.e. $P_{i}\perp\mathbf{P}^{a}$),
    • $b_{i}=\frac{\min(1,\sup_{\mathbf{y}\in\mathbb{R}^{N}}h_{i}(\mathbf{y}))}{1-\beta}$ (under general dependence).
17:  else
18:   Construct the active statistic without $X_{i}$:
@@
19:  end if
20:end for
21:Return: The set of active statistics $\{X_{i}^{\mathrm{active}}\}_{i=1}^{N}$.
The set of active statistics $\{X_{i}^{\mathrm{active}}\}_{i=1}^{N}$ produced by Algorithm 1 is designed to be broadly compatible with a wide range of downstream multiple testing procedures, a key advantage of our framework. This design allows researchers to choose the procedure that best suits the form of the statistic produced (whether a $p$-value or an $e$-value), the dependence structure in the data, and the desired power for controlling error metrics such as the False Discovery Rate (FDR, Benjamini and Hochberg, 1995).
For instance, the resulting active $p$-values can be supplied to a spectrum of methods tailored to different dependency assumptions. These range from the classic Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg, 1995), which is powerful under independence or PRDS, to the Su procedure, which provides guarantees under the PRDN assumption (Su, 2018), and the highly robust Benjamini-Yekutieli (BY) procedure (Benjamini and Yekutieli, 2001) for arbitrary dependence. Moreover, they are compatible with more advanced techniques, including adaptive procedures that estimate the null proportion $\pi_{0}$ to boost power (Storey, 2002; Storey et al., 2004) and sophisticated conditional calibration methods like the dBH procedure (Fithian and Lei, 2022).
Alternatively, when formulated as $e$-values, our statistics can be integrated with modern $e$-value-based methods, which are particularly appealing for their robustness to complex dependencies. Notable examples include the standard e-BH procedure for arbitrary dependence (Wang and Ramdas, 2022), enhanced methods that boost power via conditional calibration such as e-BH-$C$ (Lee and Ren, 2024), and unifying frameworks like the e-Closure Principle introduced by Xu et al. (2025a), which can offer uniform improvements in power and flexibility. Our framework thus serves as a flexible front-end, compatible with this entire suite of modern statistical machinery.
@@ 4 Numerical Experiments
We conduct numerical simulations to evaluate our budgeted active inference framework, comparing its statistical power and efficiency against several baselines under a fixed budget. Across all experiments we use the function $u_{i}(x)=x$ for direct signals and $u_{i}(x)=1/(x+\epsilon)$ for inverse signals as our base utility functions. Any necessary range compression to bound extreme values is automatically handled by the adaptive constraint loop built within Algorithm 1.
@@ 4.1 Competing Methods and Evaluation Metrics
We compare Algorithm  1, referred to as “Active-Default”, with the following methods:
@@
    ALL (Oracle). This non-budgeted oracle method computes the exact statistic ($E_{i}$ or $P_{i}$) for all NN hypotheses. It serves as an upper bound on statistical power at a fixed cost of NN queries.
@@
    Xu (Xu et al., 2025b). This method makes an independent probabilistic decision for each hypothesis. A Bernoulli trial $T_{i}\sim\mathrm{Bernoulli}(p_{i})$ determines whether to compute the expensive statistic, where the probability $p_{i}$ is a function of the auxiliary statistic and a hyperparameter $\beta$. For $e$-values, the query probability is $p_{i}=\max\{0,1-\beta/E_{i}^{a}\}$, and the final statistic is $E_{i}^{\mathrm{active}}=(1-T_{i})E_{i}^{a}+T_{i}(1-\beta)E_{i}$. For $p$-values, $p_{i}=\max\{0,1-\beta P_{i}^{a}\}$, and the final statistic is $P_{i}^{\mathrm{active}}=(1-T_{i})P_{i}^{a}+T_{i}(1-\beta)^{-1}P_{i}$. Crucially, because these decisions are made independently for each hypothesis, the total number of queries $\sum_{i}T_{i}$ is a random variable and is not constrained by a pre-specified global budget.
@@ 4.2 Performance with an Auxiliary Signal
In this experiment, we assess performance in a scenario where the auxiliary statistic provides a direct but unquantifiable signal about the true effect. Additional simulations are provided in Section C of the Supplement. We simulate $N=10,000$ hypotheses, each defined by a signal strength parameter $\mu_{i}$. The $i$th null hypothesis is $H_{0,i}:\mu_{i}=0$. The signal strengths $\{\mu_{i}\}_{i=1}^{N}$ are generated independently from a two-component mixture model to create a fraction $\pi$ of non-nulls:
@@
Let $\tau^{2}=2\log N$. From each primary observation $Z_{i}\sim\mathcal{N}(\mu_{i},1)$, we construct a corresponding gold-standard $e$-value and $p$-value: $E_{i}=\exp\left(\lambda Z_{i}-\frac{\lambda^{2}}{2}\right)\text{ and }P_{i}=1-\Phi(Z_{i}),$ where $\lambda=\sqrt{\log(N/\alpha)}$ as recommended in Xu et al. (2025b) and $\Phi$ is the standard normal CDF. The corresponding auxiliary statistics, which encode the signal strength $\mu_{i}$, are generated as: $E_{i}^{a}\sim\mathrm{Poisson}(1+\mu_{i})\text{ and }P_{i}^{a}\sim\mathrm{Beta}(1,1+\mu_{i}).$ The Poisson statistic serves as a direct signal for the $e$-value, while the Beta statistic provides an inverse signal for the $p$-value.
We conduct two analyses based on this setup, with a computational budget of $n_{b}=500$. First, to assess performance as a function of signal density, we vary the non-null proportion $\pi$ from 0.05 to 0.3 while holding $\beta=0.5$ fixed. Second, to examine the influence of the $\beta$, we vary it from 0.1 to 0.9 while keeping $\pi=0.1$ fixed. The target FDR level is 0.1. Given that the statistics are generated independently for each hypothesis, we employ the active $p$-value construction designed for the independent case as in (6). The results for each analysis, averaged over 100 simulations, are presented in Figures 1 and 2, respectively.